OpenStudy (anonymous):

if i gave $100 to my brother he would have twice as much as me. if he gave me $100 we would have the same amount. How much does he have now?

OpenStudy (anonymous):

So this would create two equations that you could solve as a system of equations. Let x be the amount of money you have, and y be the amount of money your brother has. So if you gave him 100, you would be losing 100 and he would be gaining 100. Also, his new balance would then be double your new balance: \[2(x-100)=y+100\] And if he gave you 100, he would be losing 100 and you would be gaining 100. Also, your new balances would then be the same:\[x+100=y-100\] Then solve them as a system of equations. First solve the 2nd equation for x:\[x=y-200\] Then substitute x in the 1st equation:\[2\left[(y-200)-100 \right]=y+100\]\[2y-600=y+100\]\[y=700\]Then plug in 700 for y in either equation and solve for x. Then you can check the two values and make sure they work as solutions:\[x=700-200\]\[x=500\]So you have 500 and he has 700

OpenStudy (anonymous):

Thanx u helped me with the same question hayeong9 wrote...

OpenStudy (anonymous):

thx heaps